ne A solid truncated cone is made of a material ofPlease explain, thanks !
Answer #1

The resistance of this cone can be found by integration.

Let us consider a thin disk of radius r at a distance x from the right hand. Then from the geometry -

(b-r)/x = (b-a)/h

Solve for radius of disk, r = (a-b)x/h + b

Now, as we know, resistance, R = rho*l/A

Where, l is the length of the conductor and A is the area of cross-section.

The contribution to the resistance from the disk having a thickness dy is

dR = rho*dx/pi*r^2

Now putting the value of r from above and integrating it from o to h

We get -

R = (rho*h)/(pi*a*b)

Putting the values given in the question -

R = (3.9*1.29)/(3.14*0.32*0.69) = 7.26 Ohm